OpenStudy (anonymous):

Hello, I need help integrating this function: e^x/ (1+e^x)^2 Its not a homework I just havent done intergal in a while and dont remember how to do it. Can someone help me.

OpenStudy (anonymous):

Make the change of variable \[u= 1+ e^x\\ du = e^x dx \]

OpenStudy (anonymous):

Don't remember how to do it.

OpenStudy (anonymous):

The integral becomes \[ \int \frac { du} { u^2}= -\frac 1 u + C = -\frac 1 {1+e^x} + C \]

OpenStudy (anonymous):

Ok, I tried it and got the same thing so that is the final answer right? no more changes to make to it? Thanks eliassaab!

OpenStudy (anonymous):

Yes, you are done if you did it as in my post above.

OpenStudy (anonymous):

hello, can someone help me to put u in terms of x please: a=parameter exp(x)=e^x a*exp(-x)*((exp(-exp(-x)))^a)=u Meant to find x in terms of u. so the form x="..."

OpenStudy (anonymous):

Hi, Anyone know how to differentiate this for example. 4^x ? a constant (4) to the power of x. Would appreciate some help.

OpenStudy (anonymous):

\[4^x \ln(4)\]

OpenStudy (anonymous):

If there are 2 elements A and B among a total of N elements and I draw a bunch of size 'n' one element after the other without replacement , what is the probability that only A is selected but not B. in terms of n and N. Anyone want to tackle this one please?